100 mL of a water sample consumed 10 mL of N/50 H

2

SO

4

during titration using


phenolphthalein as

an

indicator. Few drops of methyl orange were added to the same


solution and titration was continued. The yellow colour of the solution just turned orange


after addition of another 21 mL of the same acid solution. Ca

lculate the type and extent


of alkalinity present in the water sample.

What could be the disadvantages of using this


water sample source in boiler?



1
Expert's answer
2022-03-12T02:00:23-0500

Phenolphthalein alkalinity "=\frac{10×\frac{1}{50}×50000}{100}=100ppm"


Methylorange alkalinity "=\frac{21×\frac{1}{50}×50000}{100}=210ppm"



The alkalinity is due to CO"_3^{2-}" and HCO"_3^-"


HCO3- alkalinity "=210-100"

"=110ppm"


CO32- alkalinity "=100\u00d72"

"=200ppm"


Disadvantage

The water will form Calcium and Magnesium Carbonate in boilers.

Solid Carbonate is a bad conductor of heat, hence more electricity or fuel is needed to heat the water.

Thus wastes energy.